# Spatio-spectral diarization of meetings by combining TDOA-based segmentation and speaker embedding-based clustering

### v1

@@ 2.3 Segment-level beamforming
The segment-forming process can also result in superfluous segments that are caused by phantom positions. Before beamforming, these segments need to be identified and removed to prevent using a speaker’s own reflection as an interferer during beamforming. Reflections are characterized by the fact that its SCM shows a stronger deviation from the prototype SCM at high frequencies, which is caused by phase errors so that the corresponding binary masks are sparsely populated for higher frequencies. Therefore, the average mask activity between is compared against a threshold to determine whether the activity matches that of a speech signal. Segments containing too little activity are declared as caused by reflections and discarded.
@@ 3.3 Comparison to other systems
It can be seen that the proposed spatio-spectral system outperforms both a purely spectral and spatial approach. This shows that the proposed system effectively combines both systems’ advantages. Here, the spectral system shows stable, but lower performance due to solely using single-channel information, while the spatial model shows higher errors due to a coarser resolution and reflections in the environment.
@@
After a closer analysis, these errors could be traced back to a single loudspeaker used during the recordings exhibiting a low-pass characteristic. This loudspeaker significantly attenuates frequencies above , as can be seen in Fig. 2. Therefore, the filtering stage aimed at identifying reflections through the energy distribution of a speech signal inadvertently removes this loudspeaker’s signal before embedding extraction. The cACGMM-based mask refinement mitigates this effect and fills in gaps in the estimated speech activity, but cannot compensate for completely missed segments. With refinement, the proposed system approaches the results of the state-of-the-art SSND system [15] on LibriCSS. While SSND also employs a spatio-spectral approach, it employs a fully data-driven model to cascade diarization and separation. Due to the transformer-based multi-channel EEND network used for diarization, it requires matching training data on the corresponding microphone array geometry. Compared to it, the proposed model only requires training on VoxCeleb [36] for the embedding extractor and is agnostic to microphone placement and geometry.
@@ 3.4 Spatio-spectral diarization for changing positions
So far, all evaluations consider a static scenario with constant speaker positions. To verify the proposed, spatio-spectral pipeline’s robustness against position changes, a “semi-static” LibriCSS and LibriWASN scenario is created as follows. Two meetings are concatenated so that each overlap subset (0S – OV40) consists of long meetings, where the speaker positions change after the first half of the meeting. Here, on average five out of the eight speakers are replaced by new speakers, while three speakers change location, resulting in meetings with - different speakers but only speaker positions. This allows checking the system’s performance both in case of different sources from the same location and speakers who change their position during the meeting.

### v2

@@ 2.3 Segment-level beamforming
The segment-forming process can also result in superfluous segments that are caused by phantom positions. Before beamforming, these segments need to be identified and removed to prevent using a speaker's own reflection as an interferer during beamforming. Reflections are characterized by the fact that its SCM shows a stronger deviation from the prototype SCM at high frequencies, which is caused by phase errors so that the corresponding binary masks are sparsely populated for higher frequencies. Therefore, the average mask activity between is compared against a threshold to determine whether the activity matches that of a speech signal. Segments containing too little activity are declared as caused by reflections and discarded.
@@ 3.3 Comparison to other systems
It can be seen that the proposed spatio-spectral system outperforms both a purely spectral and spatial approach. This shows that the proposed system effectively combines both systems' advantages. Here, the spectral system shows stable, but lower performance due to solely using single-channel information, while the spatial model shows higher errors due to a coarser resolution and reflections in the environment.
@@
After a closer analysis, these errors could be traced back to a single loudspeaker used during the recordings exhibiting a low-pass characteristic. This loudspeaker significantly attenuates frequencies above , as can be seen in Fig. 2. Therefore, the filtering stage aimed at identifying reflections through the energy distribution of a speech signal inadvertently removes this loudspeaker's signal before embedding extraction. The cACGMM-based mask refinement mitigates this effect and fills in gaps in the estimated speech activity, but cannot compensate for completely missed segments. With refinement, the proposed system approaches the results of the state-of-the-art SSND system [15] on LibriCSS. While SSND also employs a spatio-spectral approach, it employs a fully data-driven model to cascade diarization and separation. Due to the transformer-based multi-channel EEND network used for diarization, it requires matching training data on the corresponding microphone array geometry. Compared to it, the proposed model only requires training on VoxCeleb [36] for the embedding extractor and is agnostic to microphone placement and geometry.
@@ 3.4 Spatio-spectral diarization for changing positions
So far, all evaluations consider a static scenario with constant speaker positions. To verify the proposed, spatio-spectral pipeline's robustness against position changes, a “semi-static” LibriCSS and LibriWASN scenario is created as follows. Two meetings are concatenated so that each overlap subset (0S – OV40) consists of long meetings, where the speaker positions change after the first half of the meeting. Here, on average five out of the eight speakers are replaced by new speakers, while three speakers change location, resulting in meetings with - different speakers but only speaker positions. This allows checking the system's performance both in case of different sources from the same location and speakers who change their position during the meeting.